OpenStudy (anonymous):

Write the equation of a line perpendicular to the line y = x - 3 and containing the point (8, -2).

OpenStudy (anonymous):

Is there a formula for this or something?

OpenStudy (anonymous):

Perpendicular lines have negative reciprocal slopes. So what is the slope of the equation you have, and what is the negative reciprocal of it?

OpenStudy (anonymous):

(That's just the first step)

OpenStudy (anonymous):

if this gradient/slope = 1 then the perpendicular gradient/slope= -1 therefor the equation = y=-x-3

OpenStudy (anonymous):

not quite, @ayesha05. y = x - 3. The slope is 1, and the negative reciprocal is -1. next, you need to take that slope and your point (8, -2) and put them in point slope form. do you know what that is, @ilmc1?

OpenStudy (anonymous):

I got it- its y=x-3 y=-x+c -2=-8+c c=6 y=-x+6

OpenStudy (anonymous):

yes thats what i was also wondering what happened to the -3

OpenStudy (anonymous):

Thanks guys c:

OpenStudy (anonymous):

you're welcome!